# Dynamical friction and measurements of the splashback radius in galaxy clusters

### v1

@@ Abstract
The splashback radius is one popular method to constrain the size of galaxy clusters. It is typically measured through the logarithmic derivative of the galaxy number density profile, since doing so is more observationally viable and computationally inexpensive compared to other methods. However, measuring the splashback radius through the galaxy number density has consistently produced smaller values of the splashback radius than those measured with dark matter density or other processes. Dynamical friction has been posited as one possible reason that splashback radii measured through galaxy number densities are reduced, since it decays the orbits of subhaloes within the halo, however, the effects of dynamical friction cannot be isolated within cosmological simulations. Here, we present idealized simulations starting with isolated galaxy clusters drawn from the IllustrisTNG cosmological simulation, where we isolate dynamical friction. We show that although dynamical friction can reduce measurements of the splashback radius, it does not have a significant effect on clusters with $M_{\mathrm{200,mean}}>10^{14}\mathrm{M_{\odot}}$, and thus cannot completely account for previously measured discrepancies.
@@ 1 Introduction
The environment of the galaxy cluster is an incredibly challenging one to understand. Dominated by dark matter and the hot, gaseous, intracluster medium, it is notoriously difficult to observe. Galaxy clusters, the most massive gravitationally bound objects in the Universe, can have masses exceeding 1015 M and may contain anywhere from hundreds to hundreds of thousands of galaxies (henceforce called cluster galaxies; see e.g. Kravtsov & Borgani, 2012, for a review). These resident cluster galaxies are typically much easier to observe than the underlying content of the cluster itself, thanks to their dominant stellar component. As such, they provide an excellent opportunity to understand the cluster environment through ensemble measurements of many galaxies.
@@
Compared to the rest of the Universe, the cluster environment is dense and hot, which is the result of strong accretion shocks, and feedback from active galactic nuclei (AGN; see McNamara & Nulsen, 2012; Kravtsov & Borgani, 2012). On average, galaxies residing in clusters do not resemble their counterpart field galaxies (galaxies that do not reside in clusters or groups): they are typically redder and more elliptical (Dressler, 1980; Hogg et al., 2004; Kauffmann et al., 2004; Blanton et al., 2005; Li et al., 2006), and have lower star formation rates (Poggianti et al., 1999; Gómez et al., 2003; O’Neil et al., 2023). Previous studies have sought to understand these differences, notably the rapid cessation of star formation (quenching) in galaxies as they fall into the cluster, identified for instance by the strength of the 4000Å break (Dressler & Shectman, 1987; Kimble et al., 1989; Kauffmann et al., 2004; Li et al., 2006; Demarco et al., 2010; Bluck et al., 2020; Kim et al., 2022). One explanation, known as “environmental quenching”, attributes the lack of star formation in cluster galaxies to a variety of processes that cause galaxies to lose their gas envelopes as they enter the cluster, such as ram pressure (seen observationally in Vollmer et al., 2004; Chung et al., 2009; Ebeling et al., 2014) and tidal stripping (Larson et al., 1980; Diemand et al., 2007).
@@
Ram pressure stripping results from interactions between the hot cluster gas and the gas in an infalling galaxy, which removes gas from the galaxy and thus lowers star formation rates (Gunn & Gott, 1972; Abadi et al., 1999; Poggianti et al., 1999), especially in the center of clusters (Wetzel et al., 2013; Borrow et al., 2023). Cluster galaxies are also impacted by tidal stripping, which is a gravitationally driven process in which mass at the edges of an infalling galaxy is pulled off. As a result, infalling galaxies lose mass, including gas. Tidal stripping has been observed in many N-body simulations (Dekel et al., 2003; Diemand et al., 2007), and like ram pressure stripping, is thought to operate most efficiently in the center of galaxy clusters (Moore et al., 1998). Subhaloes can also affect each other gravitationally in a process called harassment (Moore et al., 1998; Wetzel et al., 2013), or even merge.
@@
A foundational property of a galaxy cluster is its size. One common approach for measuring cluster size is to define a radius at which the average enclosed density is some multiple of the average density of the Universe. For instance, $R_{200,\mathrm{mean}}$ refers to the radius at which the average enclosed density is two hundred times greater than the mean density of the Universe (Peebles, 1993; Peacock, 1999). The mass enclosed by $R_{200,\mathrm{mean}}$ is accordingly referred to as $M_{200,\mathrm{mean}}$. However, this metric is defined relative to the Universe as a whole. Thus, as the density of the Universe changes over cosmic time, $R_{200,\mathrm{mean}}$ will change even if the cluster does not, a process known as pseudoevolution (Diemer et al., 2013). Consequently, the value of $M_{200,\mathrm{mean}}$ changes with time based on both the physical evolution of the cluster, and the evolution of the Universe at large.
@@
To avoid pseudoevolution, the splashback radius, or $R_{\mathrm{sp}}$, has gained popularity in recent years. The splashback radius characterizes the size of the cluster based on the radius of the first apocenter of infalling material. This material produces a “pile-up” of matter around that first turn-around point due to the reduced velocity of subhaloes at their apocenters (Adhikari et al., 2021). As the splashback radius is defined with respect to the dynamics of the cluster, it does not experience pseudoevolution.
[…]
@@
For perfectly spherical systems, this caustic and the apocenter are well-matched (Adhikari et al., 2014; Mansfield et al., 2017). However, in reality, this correlation is not perfect. Clusters are not perfectly spherical – their density profiles are greatly influenced by their merger history (Diemer & Kravtsov, 2014; Diemer, 2017; Mansfield et al., 2017; Xhakaj et al., 2020), which can reduce the strength of the caustic. Nevertheless, because of its simplicity, the differential density profile remains the primary way of measuring $R_{\mathrm{sp}}$.
Due to the long time scales involved in galaxy formation, numerical simulations have become an indispensable tool for understanding cluster evolution. Early dark matter-only simulations to investigate large-scale structure considered solely gravitational forces, and included only thousands of particles (Press & Schechter, 1974; Davis et al., 1985). Dark matter simulations are thought to be extremely precise, because they include only gravitational forces (Springel, 2010), however such simulations neglect many important processes, reducing their accuracy. The inclusion of baryonic matter introduces a tremendous challenge: simulating the numerous additional physical processes, such as hydrodynamics and radiative transfer, that this ordinary matter undergoes (Somerville & Davé, 2015; Vogelsberger et al., 2020).
Modern cosmological N-body simulations can track dark matter, as well as gas, stellar, and supermassive black holes, through time. Newer simulations include increasingly complex interactions and processes that are thought to affect galaxy evolution and the formation of large-scale structure, such as feedback from AGN and massive stars (Crain et al., 2015; Schaye et al., 2015). There are many simulations that are large enough to include galaxy clusters, which have characteristic radii on the scale of 1 Mpc. By the 2010s, projects such as Illustris and EAGLE reached simulation volumes of 106.53 and 1003 Mpc3 respectively (Vogelsberger et al., 2014; Genel et al., 2014; Crain et al., 2015; Schaye et al., 2015), while resolving down to a “particle” mass scale of $\approx 10^{6}$ M for baryonic components. Yet precisely replicating these feedback processes is persistently difficult, and mismatches with observational results remain.
Splashback radii have been measured in simulation data (Lau et al., 2015; More et al., 2016; Busch & White, 2017; O’Neil et al., 2022), as well as in numerous observational projects using data from a variety of sources with different detection techniques, such as SDSS, DES, the Subaru Telescope, and ACT (More et al., 2016; Baxter et al., 2017; Chang et al., 2018; Shin et al., 2019; Zürcher & More, 2019; Murata et al., 2020). Cluster splashback radii have been found to depend strongly on the accretion rate of the cluster, which refers to the rate at which the cluster gains mass (Diemer & Kravtsov, 2014; Shin & Diemer, 2023). Haloes with lower accretion rates possess larger splashback radii, relative to $R_{200,\mathrm{mean}}$, than those with higher mass accretion rates (More et al., 2015).
Still, much about the exact position of $R_{\mathrm{sp}}$ remains unknown. Measuring the mass density profile of a halo, which is dominated by dark matter, is challenging in observation; ideally, the galaxy number density profile would act as a useful observational substitute for the dark matter density profile. However, simulations reveal a discrepancy between the dark matter density profile and the galaxy number density profile, affecting the position of the splashback radius. This disagreement has been identified in numerous simulations and observational studies (More et al., 2016; Bahé et al., 2019; O’Neil et al., 2021; McAlpine et al., 2022).
@@
Some researchers have posited that this discrepancy is due to dynamical friction. DF affects only the subhaloes (and the associated galaxies), reducing the extent of their orbits, but not the background dark matter. Therefore, it is a possible explanation for this discrepancy between the dark matter and galaxy density profiles.
For example, using a toy model, Adhikari et al. (2016) simulated dynamical friction in a galaxy cluster system, finding that DF reduced the splashback radius, when measured through galaxy number densities, in a spherical system where all galaxies moved in perfectly radial orbits. Adhikari et al. (2016) concluded not only that dynamical friction resulted in a noticeable decrease in $R_{\mathrm{sp}}$ for subhaloes above 1% of the host halo mass, but also that the influence of DF is highly dependent on the cluster potential profile.
Bahé et al. (2019) examined the Hydrangea simulations, part of the C-EAGLE project, and argue that the impact of dynamical friction is strongest among low-mass clusters. Xhakaj et al. (2020), also using simulations, found reduced splashback radii measured with subhalo profiles, which would be consistent with the impact of dynamical friction; however, they did not investigate the source of the discrepancy. McAlpine et al. (2022) presented results from SIBELIUS-DARK, that compared the differential density profiles of the dark matter particles to those of the member galaxies. They found that using the galaxies as tracers resulted in a splashback radius approximately 10% smaller than the radius measured with dark matter particles, and state that this is consistent with dynamical friction as the cause of this discrepancy.
However, some studies argue that dynamical friction’s influence is overstated. For example, Contigiani et al. (2019) state that dynamical friction is insignificant at splashback scales. van den Bosch et al. (2016) argue that DF is only significant during the first orbital period, and only for galaxies with halo masses above 0.1 times the mass of the cluster. Dacunha et al. (2022) investigated IllustrisTNG and, while they generally attributed the decreasing splashback radii to dynamical friction, noted that DF could not explain why increases in galaxy brightness above $i<-19.4$ did not produce additional decreases in the splashback radius. A similar insensitivity to galaxy magnitude was found by More et al. (2016).
@@
Recently, O’Neil et al. (2022) find, also using IllustrisTNG, that the discrepancy between splashback radii measured through dark matter and galaxies depends on the absolute subhalo mass, rather than the mass ratio between the subhalo mass and total halo mass (as expected with the effects of DF), ultimately arguing that selection effects among galaxies are more likely the cause of the smaller-than-expected splashback radii.
@@
Observationally, Chang et al. (2018) found that high luminosity galaxies traced out smaller $R_{\mathrm{sp}}$, which is consistent with either the expected influence of dynamical friction or galaxy selection effects; however, they also found this effect to be small and within their measurement uncertainty. Murata et al. (2020) did not find consistent trends in splashback radius values with varying magnitude cuts.
@@
The issue of dynamical friction’s influence in galaxy clusters – particularly on measurements of the splashback radius – is thus a topic that is often discussed, but has not often been deeply studied. Large volume cosmological simulations involve dynamical friction as a result of N-body dynamics. This means that dynamical friction cannot be separated from the other gravitational forces. To target dynamical friction, we must run idealised simulations without the effect at all. In this paper, we seek to address this question by evolving simulated galaxy clusters with an analytical implementation of dynamical friction, allowing the role of dynamical friction to be isolated and clearly understood. The rest of the paper is structured as follows. Section 2 outlines our methods, including the implementation of dynamical friction and the simulations (Sec. 2.3). Our results are described in Section 3, and conclusions are shown in Sec. 4.
2 Methods, Observations, Simulations etc.

### v2

@@ Abstract
The splashback radius is one popular method of constraining the size of galaxy clusters, often measured through the logarithmic derivative of the galaxy number density profile. However, measuring the splashback radius through the galaxy number density has consistently produced smaller values of the splashback radius than those inferred from the underlying gravitational potential in simulations. Dynamical friction has been posited as one possible reason that splashback radii measured through galaxy number densities are reduced, since it decays the orbits of subhaloes within the halo. Dynamical friction is an emergent process, and as such, cannot be isolated or removed within N-body simulations. Here, we present simulations starting with isolated galaxy clusters drawn from the IllustrisTNG cosmological simulation, where we explicitly control dynamical friction through an idealized model. We show that although dynamical friction can reduce measurements of the splashback radius, it does not have a significant effect on clusters with $M_{\mathrm{200,mean}}>10^{14}\mathrm{M_{\odot}}$, and thus cannot account for previously measured discrepancies.
@@ 1 Introduction
The environment of the galaxy cluster is a challenging one to understand. Galaxy clusters, the most massive gravitationally bound objects in the Universe, can have masses exceeding 1015 M and may contain anywhere from hundreds to thousands of galaxies (henceforth called cluster galaxies; see e.g. Kravtsov & Borgani, 2012, for a review). Compared to the rest of the Universe, the cluster environment is dense and hot; the intracluster medium, or ICM, has densities $\sim 10^{-3}~\mathrm{cm^{-3}}$ (see e.g. Eckert et al., 2012). The deep gravitational potential well, as well as strong accretion shocks and feedback from active galactic nuclei (AGN; see McNamara & Nulsen, 2012; Kravtsov & Borgani, 2012) heat the gas to temperatures $\sim 10^{7}-10^{8}$ K, producing X-rays via thermal bremsstrahlung (e.g. Leccardi & Molendi, 2008; Chen et al., 2024). Clusters can thus be detected through this X-ray emission (see e.g. Böhringer et al., 2001; Pacaud et al., 2016), as well as through the Sunyaev-Zeldovich effect (SZ effect; see e.g. Bleem et al., 2015; Planck Collaboration et al., 2016b).
@@
Clusters are also selected and studied through overdensities in their resident galaxies, in the optical (see e.g. Kepner et al., 1999; Rykoff et al., 2014; Sarron et al., 2018; Huang et al., 2021) and near-infrared (see e.g. Wylezalek et al., 2013; Rettura et al., 2014; Gully et al., 2024, at $z>1.3$). On average, galaxies residing in clusters do not resemble their counterpart field galaxies (galaxies that do not reside in clusters or groups): they are typically redder and more elliptical (Dressler, 1980; Hogg et al., 2004; Kauffmann et al., 2004; Blanton et al., 2005; Li et al., 2006), and have lower star formation rates (Poggianti et al., 1999; Gómez et al., 2003; O’Neil et al., 2023). Previous studies have sought to understand these differences, notably the rapid cessation of star formation (quenching) in galaxies as they fall into the cluster, identified for instance by the strength of the 4000 Å break (Dressler & Shectman, 1987; Kimble et al., 1989; Kauffmann et al., 2004; Li et al., 2006; Demarco et al., 2010; Bluck et al., 2020; Kim et al., 2022).
@@
One explanation, known as “environmental quenching”, attributes the lack of star formation in cluster galaxies to a stripping processes that cause galaxies to lose their gas envelopes as they enter the cluster. Ram pressure stripping results from interactions between the hot cluster gas and the gas in an infalling galaxy, which removes gas from the galaxy and thus lowers star formation rates (Gunn & Gott, 1972; Abadi et al., 1999; Poggianti et al., 1999; Vollmer et al., 2004; Chung et al., 2009; Ebeling et al., 2014), especially in the center of clusters (Wetzel et al., 2013; Borrow et al., 2023). Cluster galaxies are also impacted by tidal stripping, which is a gravitationally driven process in which mass at the edges of an infalling galaxy is pulled off. As a result, infalling galaxies lose mass, including gas. Tidal stripping has been observed in many N-body simulations (Dekel et al., 2003; Diemand et al., 2007), and like ram pressure stripping, is thought to operate most efficiently in the center of galaxy clusters (Larson et al., 1980; Moore et al., 1998). Subhaloes can also affect each other gravitationally in a process called harassment (Moore et al., 1998; Wetzel et al., 2013), or even merge. Additionally, infalling galaxies can be “pre-processed” by smaller galaxy groups as they fall into the cluster along filaments (Fujita, 2004; Sarron et al., 2019; Lee et al., 2022).
@@
A foundational property of a galaxy cluster is its size, and a variety of metrics are used to measure cluster size, such as the virial radius $R_{\mathrm{vir}}$ assuming a spherical collapse model (Bryan & Norman, 1998). Based on the spherical collapse model, a related approach for measuring cluster size is to define a radius at which the average enclosed density is some multiple of a reference density (see Cole & Lacey, 1996; Diemer et al., 2013). For instance, $R_{200,\mathrm{mean}}$ refers to the radius at which the average enclosed density is two hundred times greater than the mean density of the Universe (Peebles, 1993; Peacock, 1999), which in simulations can be easily calculated for the box volume. The mass enclosed by $R_{200,\mathrm{mean}}$ is accordingly referred to as $M_{200,\mathrm{mean}}$. (Other common analogous radii include $R_{200,\mathrm{crit}}$, $R_{500,\mathrm{mean}}$, and $R_{500,\mathrm{crit}}$, which are more useful for studying the inner regions of haloes, as discussed by Diemer et al., 2013). However, since this metric is defined relative to the Universe as a whole, thus, as the density of the Universe changes over cosmic time, $R_{200,\mathrm{mean}}$ will change along with the Universe to maintain a constant density ratio, a process known as pseudoevolution (Diemer et al., 2013), regardless of the accretion of material. This phenomenon has been measured in numerous N-body simulations where the virial mass increases without a corresponding level of accretion (Prada et al., 2006; Diemand et al., 2007; Zemp, 2014; Shirasaki, 2019; Wang et al., 2020). In particular, Wetzel & Nagai (2015) found that the dark matter density profile is strongly affected by pseudoevolution. The consequence of pseudoevolution is that the value of $M_{200,\mathrm{mean}}$ changes with time based on both the physical evolution of the cluster, and the evolution of the Universe at large.
@@
To avoid pseudoevolution and more accurately trace halo mass accretion histories, other methods of measuring halo boundaries have emerged, for example, the turnaround radius or the depletion radius (see e.g. Pavlidou & Tomaras, 2014; Fong & Han, 2021, respectively). In particular, the splashback radius, or $R_{\mathrm{sp}}$, has gained attention in recent years (Adhikari et al., 2014). The splashback radius characterizes the size of the cluster based on the radius of the first apocenter of infalling material, thus probing the mass accretion history (Diemer & Kravtsov, 2014; More et al., 2015). This material produces a “pile-up” of matter around that first turn-around point due to the reduced velocity of subhaloes at their apocenters (Adhikari et al., 2021). As the splashback radius is defined with respect to the dynamics of the cluster, it does not experience pseudoevolution but evolves between $\sim$ 1-2 $R_{\mathrm{200,mean}}$, varying with mass, redshift, and mass accretion rate (see e.g. Diemer, 2017; O’Neil et al., 2021; Towler et al., 2024).
@@
The dynamical splashback feature is associated with a caustic in the density profile. In particular, the logarithmic density profile ($\rho/dlog$$r$) exhibits a dip in the profile at this region. The local minimum of that feature marks an excellent proxy for $R_{\mathrm{sp}}$ (Diemer & Kravtsov, 2014). Due to observational and computational challenges in finding the apocenters of numerous cluster galaxies or particles, the density profile is therefore a commonly used proxy for finding the splashback radius (see e.g. More et al., 2015; Chang et al., 2018; Shin et al., 2019; Murata et al., 2020; O’Neil et al., 2021). For perfectly spherical systems, this caustic and the apocenter are well-matched (Adhikari et al., 2014; Mansfield et al., 2017). However, in reality, this correlation is not perfect. Clusters are not perfectly spherical, which can reduce the strength of the caustic (Diemer & Kravtsov, 2014; Diemer, 2017; Mansfield et al., 2017; Xhakaj et al., 2020). Nevertheless, because of its simplicity, the differential density profile remains the primary way of measuring $R_{\mathrm{sp}}$. In particular, the galaxy number density profile is commonly used because galaxy overdensities are a common tracer of clusters.
Splashback radii have been measured in optically-selected clusters from weak lensing mass measurements (e.g. Chang et al., 2018; Fong et al., 2022; Giocoli et al., 2024; Xu et al., 2024) and through galaxy number density (e.g. More et al., 2016; Baxter et al., 2017; Shin et al., 2019; Murata et al., 2020; Bianconi et al., 2021; Kopylova & Kopylov, 2022; Contigiani et al., 2023), as well as in X-ray selected clusters, similarly with weak lensing (Umetsu & Diemer, 2017; Contigiani et al., 2019) and number density of galaxies (Bianconi et al., 2021; Rana et al., 2023). SZ-selected clusters have also been studied through galaxy density profiles (Shin et al., 2019; Zürcher & More, 2019; Adhikari et al., 2021; Shin et al., 2021), weak lensing (Shin et al., 2021), and even gas pressure profiles (Anbajagane et al., 2022). Gonzalez et al. (2021) even detected the splashback radius in intracluster light, perhaps also the first detection of the splashback radius for an individual cluster, rather than stacked haloes.
@@
In simulation, splashback radii have been detected dynamically through particle orbit tracing (e.g. finding the particle apocenters Diemer et al., 2017; García et al., 2023; Shin & Diemer, 2023), as well as through dark matter density profiles in N-body and hydrodynamic simulations (Lau et al., 2015; Aung et al., 2021; Deason et al., 2021; O’Neil et al., 2021; McAlpine et al., 2022; Haggar et al., 2024; Lebeau et al., 2024). Baryonic contributions have also been studied (e.g. Zhang et al., 2023; Towler et al., 2024) in gas (Lau et al., 2015; O’Neil et al., 2021; Lebeau et al., 2024) and subhalo number densities (see e.g. O’Neil et al., 2021; Dacunha et al., 2022; O’Neil et al., 2022; Pizzardo et al., 2024, all using IllustrisTNG data).
Cluster splashback radii have been found to depend strongly on the accretion rate of the cluster, which refers to the rate at which the cluster gains mass (Diemer & Kravtsov, 2014; Shin & Diemer, 2023). Haloes with lower accretion rates possess larger splashback radii, relative to $R_{200,\mathrm{mean}}$, than those with higher mass accretion rates (More et al., 2015).
Much about the exact relationship between the differential density profile and the splashback radius, as defined from dynamics, remains unknown. Further, even the systematics relating differential density profiles measured from different proxies are poorly understood. Measuring the mass density profile of a halo, which is dominated by dark matter, is challenging in observation; ideally, the galaxy number density profile would act as a useful observational substitute for the dark matter density profile. However, simulations reveal a discrepancy between the dark matter density profile and the galaxy number density profile, affecting the position of the splashback radius. This disagreement has been identified in numerous simulations and observational studies (see e.g. More et al., 2016; Bahé et al., 2019; O’Neil et al., 2021; McAlpine et al., 2022).
@@
Some researchers have posited that this discrepancy is due to dynamical friction, largely using simulations (see e.g. Bahé et al., 2019; Xhakaj et al., 2020; McAlpine et al., 2022). Dynamical friction affects primarily the subhaloes (and the associated galaxies), reducing the extent of their orbits, but not the background dark matter, and consequently reducing the splashback radius measurement from galaxy number density profiles. Using a toy model, Adhikari et al. (2016) simulated dynamical friction in a galaxy cluster system, finding that DF reduced the splashback radius, when measured through galaxy number densities, in a spherical system where all galaxies moved in perfectly radial orbits.
@@
However, some studies argue that dynamical friction’s influence is overstated, for example that it is insignificant at splashback scales (Contigiani et al., 2019) or affects only a few very massive subhaloes (van den Bosch et al., 2016). Insensitivities to galaxy magnitude in splashback radius measurements (e.g. More et al., 2016; Murata et al., 2020) also challenge the expected role of DF, with Dacunha et al. (2022) noting that DF cannot explain why increases in galaxy brightness above i<19.4i<-19.4 do not produce additional decreases in the splashback radius. Recently, O’Neil et al. (2022) found, using IllustrisTNG, that the discrepancy between splashback radii measured through dark matter and galaxies depends on the absolute subhalo mass, rather than the mass ratio between the subhalo mass and total halo mass (as expected with the effects of DF), ultimately arguing that selection effects among galaxies are more likely the cause of the smaller-than-expected splashback radii.
@@
Results from observational studies are also mixed: Chang et al. (2018) found that high luminosity galaxies traced out smaller $R_{\mathrm{sp}}$, which is consistent with either the expected influence of dynamical friction or galaxy selection effects; however, they also found this effect to be small and within their measurement uncertainty. Murata et al. (2020) did not find consistent trends in splashback radius values with varying magnitude cuts. Measurements of the splashback radius through galaxy number density, with SZ- or X-ray-selected clusters have been consistent with those expected from the dark matter profile (see e.g., Zürcher & More, 2019; Shin et al., 2019; Adhikari et al., 2021; Shin et al., 2021), indicating that optical selection effects might significantly contribute to the discrepancy (see also discussion in Murata et al., 2020).
@@
The issue of dynamical friction’s influence in galaxy clusters – particularly on measurements of the splashback radius – is thus a topic that is often discussed, but has not been clearly isolated. Large volume cosmological simulations naturally solve for (potentially inaccurate, see van den Bosch & Ogiya, 2018) dynamical friction as a result of N-body dynamics. This means that dynamical friction cannot be separated from the other gravitational forces, and as such there is an extremely limited body of work studying its effects. To target DF, we must run idealised simulations that explicitly model dynamical friction such that it can be removed for comparison purposes.
In this paper, we seek to address this question by evolving simulated galaxy clusters with an analytical implementation of dynamical friction, allowing the role of dynamical friction to be isolated and clearly understood. By using a realistic population of galaxies extracted from the IllustrisTNG simulation suite, we aim to, for the first time, quantify the potential scale of the impact of DF on the splashback radius of galaxy clusters as measured using the galaxy number density profile. Our overarching aim is not to provide specific numerical values for comparisons to observations (as to perform this specific work we must make a number of key assumptions); rather this paper seeks to take the first steps in quantifying whether or not dynamical friction can impact these measurements, and for what range of galaxy and cluster masses. The rest of the paper is structured as follows. Section 2 outlines our methods, including the implementation of dynamical friction and the simulations (Sec. 2.3). Our results are described in Section 3, with discussion and conclusions shown in Secs. 4 and 5.
2 Methods